OpenStudy (trojanpoem):

Calculus : Inverse functions: Prove that: sin( x+ tan^-1(cotx)) = 1

OpenStudy (superdavesuper):

its more a trigo prob than calculus.. u can use the eqn cot(x) = tan(pi/2 - x) to simplify tan-1(cot(x))

OpenStudy (trojanpoem):

I found another solution , by assuming that tan inverse of cotx = y tany = cotx , drawing the triangle and sin(x + y) = sinx cos y + siny cos x = sin^2x + cos^2y ( from the triangle) = 1 But I am interested in knowing , how you would solve it.

OpenStudy (superdavesuper):

mine would be tan-1(cotx)=tan-1(tan(pi/2-x))=pi/2 - x the rest would follow but ur way works just fine too :)

OpenStudy (trojanpoem):

Ok, I got it sin(x + tan^-1 tan(90 - x)) = 90

OpenStudy (superdavesuper):

Correct!

OpenStudy (trojanpoem):

I believe , your method is way better.

OpenStudy (superdavesuper):

Both works :)

OpenStudy (trojanpoem):

Thanks.